# Validity of the Fröhlich model for a mobile impurity in a Bose-Einstein condensate

### v2

@@ III.2 Renormalization of the Fröhlich Hamiltonian
At this stage, ignoring cubic and quartic contributions, the resulting Hamiltonian is unitarily equivalent to $\hat{H}_{\mathrm{Bog}}$ cut off at momentum $N^{\tau}$, via a Weyl transformation similar to $U_{W}^{\dagger}$.
The renormalization at the order $N^{\tau}$ its ultraviolet divergence is performed again via a Bogoliubov and a Weyl transformation, which are however not perturbations of the identity.
@@
(11)$\mapsto\frac{2\pi\sqrt{\alpha}}{\mu}\sum_{0<k<N^{\tau}}\sqrt{\frac{k^{2}}{2m\omega_{k}}}e^{-i{\bm{X}}\cdot{\bm{k}}}b_{\bm{k}}^{\dagger}+\mathrm{h.c.}\,.$
@@
We can now finish regularizing the linear terms (11) with another Weyl transform, taking into account the correct dispersion relation $\omega$ for low-energy phonon-impurity scattering. Its action is given by
@@
This extracts a scalar of order $N^{\tau}$ which combines with the one created by $U_{W}$, yielding again a contribution of order one. It also transforms the linear terms (11) to more regular quadratic terms, completing the ones in (12). Calling $U$ the product of all the unitary transformations and taking $w({\bm{k}})$ equal to $2\pi\sqrt{\alpha}k/(\mu(\frac{1}{2}k^{2}+\omega_{k})\sqrt{2m\omega_{k}})$ for $k\leq N^{\tau}$, and $\sqrt{N}\varphi_{\mathrm{IB}}$ for $k>N^{\tau}$, we obtain up to small errors
@@
$\hat{K}=\tfrac{1}{2}\hat{P}^{2}+\sum_{{\bm{k}}}\omega_{k}b^{\dagger}_{\bm{k}}b_{\bm{k}}+\sum_{{\bm{k}},{\bm{\ell}}\neq 0}\hat{V}_{\mathrm{IB}}({\bm{k}})e^{-i{\bm{X}}\cdot{\bm{k}}}b^{\dagger}_{{\bm{k}}+{\bm{\ell}}}b_{\bm{\ell}},\$
@@
$G=$ $-\frac{1}{2\hat{K}}\sum_{k,\ell\neq 0}w({\bm{k}})({\bm{k}}\cdot{\bm{\ell}})w({\bm{\ell}})e^{-i{\bm{X}}\cdot({\bm{k}}+{\bm{\ell}})}b^{\dagger}_{{\bm{k}}}b^{\dagger}_{{\bm{\ell}}},$
@@
$(1-G^{\dagger})\hat{K}(1-G)-G^{\dagger}\hat{K}G+\hat{R},$
which eliminates the problematic interaction terms. The new term $G^{\dagger}\hat{K}G$ contains the divergent self energy. Putting this term into normal order, we find the scalar
$\langle\varnothing|\hat{G}^{\dagger}\hat{K}^{-1}\hat{G}|\varnothing\rangle\approx\frac{1}{2}\sum_{{\bm{k}},{\bm{\ell}}\neq 0}\frac{w({\bm{k}})^{2}w({\bm{\ell}})^{2}({\bm{k}}\cdot{\bm{\ell}})^{2}}{\frac{1}{2}|{\bm{k}}+{\bm{\ell}}|^{2}+\omega_{k}+\omega_{\ell}}$
plus corrections of order one due to $\hat{V}_{\mathrm{IB}}({\bm{k}})$. Evaluating the sum for large $N$ yields the $\log N$-term from Theorem 1. One easily confirms that $\langle\varnothing|\hat{R}^{\dagger}\hat{K}^{-1}\hat{R}|\varnothing\rangle$ is of order one, i.e, $\hat{R}$ is not singular. The normal ordered term $\mathbin{:}G^{\dagger}\hat{K}G\mathbin{:}$ no longer suffers from singularities and gives a finite operator also for $N=\infty$ (since it contains an additional factor of $\hat{K}^{-1}$). This is also true for $\hat{R}$. We thus have a limiting operator for $N\to\infty$ that is unitarily equivalent to the renormalized Bogoliubov-Fröhlich Hamiltonian via the inverse of the Weyl transformation (14). Note that the term with $\hat{V}_{\mathrm{IB}}({\bm{k}})$ disappears from $\hat{K}$ in the limit. This term gives a global energy shift order one, but does not change the differences of eigenvalues.
The proof of Theorem 1 relies on min-max formulas which require bounds on various error terms accumulated in the calculations. As we have explained, these are small relative to the kinetic energies, and the positive interaction terms (6h), (6f). They may thus be neglected for a lower bound on the energy. An upper bound is obtained by constructing trial states using the eigenvectors of the Bogoliubov-Fröhlich Hamiltonian and the unitary transformations.
@@ IV Conclusion
We have obtained an expansion of the energy for a dilute Bose gas coupled to an impurity with effective coupling $\alpha\sim 1$. By analyzing precisely the renormalization of the model parameters due to the emergence of a point-potential, we have shown the presence of a correction due to an effective three-body interaction. We have also established that the excitation spectrum is accurately described by the Bogoliubov-Fröhlich Hamiltonian.

### v3

@@ III.2 Renormalization of the Fröhlich Hamiltonian
At this stage, ignoring cubic and quartic contributions, the resulting Hamiltonian is unitarily equivalent to $\hat{H}_{\mathrm{Bog}}$ cut off at momentum $N^{\tau}$, via a Weyl transformation similar to $U_{W}^{\dagger}$. This operator is still ultraviolet divergent when we remove the cutoff by taking $N\to\infty$ . The renormalization of the energy at order $N^{\tau}$ is performed again via a Bogoliubov and a Weyl transformation, which are however not close to the identity.
@@
[…]
@@
We can now finish regularizing the linear terms (13) with another Weyl transform, taking into account the correct dispersion relation $\omega$ for low energy phonon-impurity scattering. Its action is given by
@@
This extracts a scalar of order $N^{\tau}$ which combines with the one created by $U_{W}$, yielding again a contribution of order one. It also transforms the linear terms (13) to more regular quadratic terms, completing the ones in (14). Calling $U$ the product of all the unitary transformations and defining $w({\bm{k}})$ to be equal to $2\pi\sqrt{\alpha}k/(\mu(\frac{1}{2}k^{2}+\omega_{k})\sqrt{2m\omega_{k}})$ for $k\leq N^{\tau}$, and $\sqrt{N}\varphi_{\mathrm{IB}}$ for $k>N^{\tau}$, we obtain up to small errors
@@
$\hat{K}=\tfrac{1}{2}\hat{P}^{2}+\sum_{{\bm{k}}}\omega_{k}b^{\dagger}_{\bm{k}}b_{\bm{k}}+\sum_{{\bm{k}},{\bm{\ell}}\neq 0}\hat{V}_{\mathrm{IB}}({\bm{k}})e^{-i{\bm{X}}\cdot{\bm{k}}}b^{\dagger}_{{\bm{k}}+{\bm{\ell}}}b_{\bm{\ell}},\\$
@@
$\hat{G}$= $-\frac{1}{2\hat{K}}\sum_{k,\ell\neq 0}w({\bm{k}})({\bm{k}}\cdot{\bm{\ell}})w({\bm{\ell}})e^{-i{\bm{X}}\cdot({\bm{k}}+{\bm{\ell}})}b^{\dagger}_{{\bm{k}}}b^{\dagger}_{{\bm{\ell}}},$
@@
$(1-\hat{G}^{\dagger})\hat{K}(1-\hat{G})-\hat{G}^{\dagger}\hat{K}\hat{G}+\hat{R}$,
which eliminates the problematic interaction terms. The new term $\hat{G}^{\dagger}\hat{K}G$ contains the divergent self energy. Putting this term into normal order, we find the vacuum expectation
$\langle\varnothing|\hat{G}^{\dagger}\hat{K}\hat{G}|\varnothing\rangle\approx\frac{1}{2}\sum_{{\bm{k}},{\bm{\ell}}\neq 0}\frac{w({\bm{k}})^{2}w({\bm{\ell}})^{2}({\bm{k}}\cdot{\bm{\ell}})^{2}}{\frac{1}{2}|{\bm{k}}+{\bm{\ell}}|^{2}+\omega_{k}+\omega_{\ell}}$
plus corrections of order one due to $\hat{V}_{\mathrm{IB}}({\bm{k}})$. Evaluating the sum for large $N$ yields the $\log N$-term from (4). The normal ordered term $\mathbin{:}\hat{G}^{\dagger}\hat{K}\hat{G}\mathbin{:}$ no longer suffers from singularities and gives a finite operator also for $N=\infty$ (since it contains an additional factor of $\hat{K}^{-1}$). This is also true for $\hat{R}$. We thus have a limiting operator for $N\to\infty$ that is unitarily equivalent to the renormalized Bogoliubov-Fröhlich Hamiltonian via the inverse of the Weyl transformation (16). Note that the term with $\hat{V}_{\mathrm{IB}}({\bm{k}})$ disappears from $\hat{K}$ in the limit. This term gives a global energy shift of order one, but does not change the differences of eigenvalues.
The proof of the eigenvalue approximation (5) relies on min-max formulas, which require bounds on various error terms accumulated in the calculations. As we have explained, these are small relative to the kinetic energies, and the positive interaction terms (8h), (8f). They may thus be neglected for a lower bound on the energy. An upper bound is obtained by constructing trial states using the eigenstates of the Bogoliubov-Fröhlich Hamiltonian and the unitary transformations.
@@ IV Conclusion
We have obtained an expansion of the energy for a dilute Bose gas coupled to an impurity with effective coupling $\alpha\sim 1$. By analyzing precisely the renormalization of the model parameters due to the emergence of a point-interaction, we have shown the presence of a correction due to an effective three-body interaction. We have also established that the excitation spectrum is accurately described by the Bogoliubov-Fröhlich Hamiltonian.